Solve.
4x^4-25x^2+36=0

If there is more than one solution, separate them with commas.

Answer by ewatrrr(24785)


Hi
4x^4-25x^2+36=0 |substition showed x = 2 is a root
(x-2)(4^3+8x^2-9x-18) |found by dividing 4x^4-25x^2+36 by (x-2)
(x-2)(4^3+8x^2-9x-18) |substitiution showed x = -2 is a root for 4^3+8x^2-9x-18
(x-2)(x+2)(4x^2-9) |found by dividing 4^3+8x^2-9x-18 by (x+2)
factoring
(x-2)(x+2)(2x-3)(2x+3)=0
x = {-2,-2,3/2,-3/2}